# CORP: A Multi-Modal Dataset for Campus-Oriented Roadside Perception Tasks

### v2

@@ 1 Introduction
Roadside perception technology (RP) is a key component in the intelligent transportation and autonomous driving domains, enabling features such as traffic flow management[1], decision support for autonomous vehicles[2], and campus surveillance and security[3], where the pursuit of reliable perception methods and high-quality datasets remain enduring topics.
@@
On a university campus, we have tested some open-sourced state-of-the-art algorithms in the roadside perception field such as BEVHeight[9], BEVDepth[10] and BEVFormer[11] with DAIR-V2X-I[4] and corresponding pre-trained models. When applied to campus roads characterized by complex living scenes and natural elements, these algorithms demonstrated unsatisfactory performance, failing to effectively distinguish between various objects and backgrounds, as illustrated in Fig. 1. This phenomenon indicates that, despite their outstanding performance on urban traffic datasets like DAIR-V2X[4] and Rope3D[5], these algorithms still need improvement in terms of generalization capability and robustness when faced with more diverse environments.
@@
The highlighted observation points towards a research frontier in which residential areas akin to university campuses provide distinctive research vantages and real-world deployment scenarios, thereby broadening the scope of roadside perception technology development beyond the traditional confines of urban streets. Motivated by this, we have innovatively constructed a roadside perception dataset named CORP (Campus-Oriented Roadside Perception), which is the first large-scale, multi-modal roadside dataset specifically designed for campus and large public spaces. Within the CORP dataset, in response to the unconstrained nature of the campus environment, multiple cameras and lidars with diverse specifications have been employed and installed at varying heights and locations on roadside poles across the campus, capturing rich scene information from various perspectives. Table 1 shows that according to the number of sensors and the complexity of scenes, we have partitioned the CORP dataset into three regions, namely Region A, Region B, and Region C, as shown in Fig. 2.
@@ 2.1 Roadside Perception Datasets
Datasets exist for perception tasks in campus settings. The ShanghaiTech[12] dataset is designed for anomaly detection with images, which contains videos taken from multiple scenes in a university in Shanghai, and hundreds of footages are labeled with various types of abnormal events. The NWPU Campus[13] is built for the same task on a larger scale. The Usyd Campus[14] dataset provides a multi-modal collection over and around the University of Sydney campus for a continuous period of 1.5 years, focusing on vehicle-side perception. The Campus3D[15] dataset was generated through photogrammetric processing of UAV images taken over the National University of Singapore (NUS), which provides a dense point cloud of a campus tailored for point segmentation tasks. A comparison between CORP and the aforementioned datasets can be found in Table 2.
@@ 3.1 Sensor Configuration
In the CORP dataset, the data frames were meticulously accumulated within a university campus, incorporating high-resolution imagery from 18 traffic surveillance cameras and spatial information captured by an ensemble of 9 automotive-grade LiDAR sensors. Each of these sensors was strategically installed atop intelligent poles at varying heights ranging from 2.5 to 6.5 meters above the ground, thereby ensuring comprehensive and diversified coverage for robust roadside perception tasks and presenting new challenges. An overview of the point cloud pattern and the position of cameras is shown in Fig. 2 with a real-world picture of the pole. The range of the whole area extends about 700 meters in the west-east direction and has a north-south length of 350 meters. We divide the area into three regions, labeled A, B, and C, based on the detection range, and color-coded with blue, red, and black, respectively. We have configured different sensor layout schemes for different regions, and the specific number and distribution details of the sensors in the layout are shown in Table 1.
@@
We have constructed four types of coordinate systems for the CORP dataset, of which two are used for LiDAR sensors, one for cameras, and the last one for the global map, as illustrated in Fig. 3. For LiDARs, the LiDAR-ego reference frame originates from the sensor itself with the $x$-axis pointing forward and $z$-axis upward. For the convenience of researchers and to safeguard data privacy, we have constructed an additional virtual LiDAR coordinate system, termed as LiDAR-base, transforms the origin of LiDAR-ego to the ground contact point of its host pole to comply its $x$-$o$-$y$ plane with the local ground surface. For cameras, the origin of the camera-ego coordinate system overlaps the optical center of the sensor with the $z$-axis pointing forward along the camera’s view and $x$-axis pointing leftward accordingly. The LiDAR and camera coordinate systems are employed to represent objects in the sensor’s field of view, while the map coordinates unify the objects into a global picture. The map coordinate system has an origin with known GPS, its $x$-axis coincides with the east and the $y$-axis with the north.
@@
Intrinsic parameters of a camera define the projection between the incoming rays and image pixels. For CORP, we use chessboards as explicit targets and employ MATLAB Camera Calibrator [50] following the method outlined by Zhang et al.[51] to determine the intrinsic calibration. This procedure yields a set of quantities including $[f_{x},f_{y},c_{x},c_{y},k_{1},k_{2},p_{1},p_{2},k_{3}]$, where $f_{x}$ and $f_{y}$ denote the focal lengths, while $c_{x}$ and $c_{y}$ represent the principal points. Additionally, the parameters $k_{1}$, $k_{2}$, and $k_{3}$ pertain to radial distortion, and $p_{1}$ and $p_{2}$ are related to tangential distortion.
@@ 3.2 Data Acquisition and Annotation
Data Labeling and Visualization. Human annotators are employed for joint labeling of the 2D and 3D targets in the dataset using the format in DAIR-V2X, with additional information such as target ID and motion status (movingv.s. static) added. To generate pixel segmentation masks for the moving instances in mages, the foundation vision model SAM[52] is applied with the integration of human intervention and manual guidance. Visualization of data samples is shown in Fig. 4. The images and point clouds can also be viewed with their annotations from SUSTechPoints[53] with our development toolkit.
@@ 3.4 Features and statistics
To get an overview of how targets are distributed in space, we aligned all identified 3D objects into their own LiDAR-base coordinate system, each of which has $x$-axis pointing eastward and $y$-axis towards the north, and visualized them collectively as shown in Fig. 5(a), where warmer color indicates a higher concentration of targets. Also, it is not surprising that the spatial distribution of targets exhibits a discernible trend toward specific directions as also demonstrated in their orientation profile in Fig. 5(c), since the mobility of targets is partially subjected to the road network in the environment. The allocation of targets indicates that our dataset encompasses plenty of road arrangements. Fig. 5(b) showed the distribution profile for the distance between targets and their center LiDARs under BEV. The portion of targets decreases as the distance increases since point clouds are sparser there, but most of the targets appear relatively evenly within 100 meters away from the center.
@@
To glance at how targets distribute differently in CORP than in urban arterial streets, we analyzed the target densities of CORP, DAIR-V2X-i[4] and Rope3D[5]. As shown in Fig. 6(a) and (c). CORP displays a sparsity of per-frame target numbers in both 2D and 3D cases, about half of the frames contain less than 15 targets while the other half has up to 50 targets per frame, which is not the case in its urban counterparts. However, the sparsity of targets does not imply simplicity of perception tasks, since accurately and efficiently localizing them from the background requires extra engineering for the model. In addition,
Although the volume of the object class is less than that of DAIR-V2X and A9, we found no significant imbalance in the number of targets, indicating a relatively minor disparity over categories, thus diminishing the risks of label inequity. Note that, Rope3D has no 2D bounding box annotations and therefore is excluded in Fig. 6(c) and (d).
@@ 4.1 Baseline Methods
For LiDAR-based baseline, we trained PointPillars with a detection range customized to [-102.4, -102.4,-2.0, 102.4, 102.4, 4.0] and the size of the voxel grid as [0.16, 0.16] along $x$ and $y$ axes. All point clouds are transformed with annotations to the LiDAR-base coordinate system mentioned in Section 3. For camera-based detection, we compared the performance of BEVHeight and ImVoxelNet. In addition, we trained BEVFusion following [43] but confined to perception sites with both camera and LiDAR installed. We adapt AB3D to the detection results as our tracking baseline.
@@ 4.2 Results
For 2D detection with CORP, we compared the performance of broadly acknowledged YOLO-V5 and its V8 counterparts. As shown in Table 4, the two methods demonstrate an obvious gap in overall precision, recall and AP scores. Despite the higher precision for YOLO-V5, V8 wins in terms of $AP_{50-95}$, implying a better generalization ability. However, challenges persist in accurately detecting distant targets, particularly when they are occluded by foreground elements or under suboptimal light conditions as shown in Fig. 7(a).
@@
For the evaluation of P3D, we compared it with the traditional yet industrially-recognized IPM method[61] in terms of prediction error in the distance from the camera to the target. For a fair comparison, images from 18 cameras in CORP are examined within their corresponding region, categorized into A, B and C based on the detection range. An experiment vehicle equipped with RTK GPS is driven through the FOV of each camera slowly, then the distance from the vehicle to the camera is predicted by P3D and IPM independently. As shown in Fig. 8, although for region A, the prediction error of P3D is larger in the middle but the overall error is less than 2 meters for that interval. Also we observed that P3D consistently outperformed IPM across all scenarios featuring diverse ground surfaces, camera height and pose angles, thereby establishing itself as a robust baseline for object-level distance estimation tasks. Again, here we focus on non-learning distance estimation methods.
@@ 5 Conclusion
In roadside scenarios, the camera sensor is usually positioned at a certain height $H_{c}$ above a local ground plane. We define the camera and the road coordinate systems as shown in Figure Fig. 9. By employing P3D, the objective we aspire to achieve is to lift the pertinent 2D objects into real-world environment by the reconstruction of their 3D coordinates, in a non-learning fashion.
@@
To describe the ground plane in the camera’s frame, we use the transformation matrix $T_{road2cam}\in R^{4\times 4}$ to establish the relationship between the road and the camera coordinates. Any 3D point in the road system $[x,y,z]_{road}^{T}$ can be transformed to the image plane to obtain their pixel coordinates $[u,v]^{T}$ via the following two steps:
@@
where $Q_{cam}$ is the coordinates of a point in the camera system, $n=\begin{bmatrix}r_{21}&r_{22}&r_{23}\end{bmatrix}^{T}$ is the normal vector to the plane and $d$ is the intercept parameter, which is $-H_{c}$ in the considered case. When it comes to a different plane, $d$, $\alpha$ and $\gamma$ should all be accounted for from that plane. By solving the above equations for $z$ in the case where $\gamma$ is close to 0, this depth $z$ can be approximated as:
@@
Once $z$ of a target in an image is obtained with its pixel coordinates $[u,v]^{T}$, we can further get its Cartesian coordinates $[x,y,z]_{cam}^{T}$ in the camera system by the following:
@@
Worth mentioning, the vanishing point of a camera, defined by the intersection of two mutually parallel lines in $o$-$x$-$y$-$z$ coordinate system, is considered to have infinite depth, i.e.the dominator in Eq. 7 on the right side is zero, therefore we have:
@@
For the completeness of description, the overview of our Pseudo-3D (P3D) method for target distance estimation is shown in Fig. 10. The processing pipeline of P3D is the following: (i) with a given image, perform line feature detection for parallel lines, which in most cases are lane markings, then find their intersections as the vanishing point. (ii) using the vanishing point to obtain the pitch angle of a camera based on Eq. 9. (iii) with the given pixel coordinates $[u,v]^{T}$ of a target, compute the estimated depth $z$ using Eq. 7 and $x$, $y$ using Eq. 8.
@@
Once the intrinsic parameters, pose angles and the height of camera sensors are measured, we can lift a 2D target in an image to a 3D point in the camera coordinate system by following the closed-form Eq. 7 and Eq. 8 with no computational cost, given an image-based 2D detector employed beforehand to produce the bounding boxes of interested targets. Subsequently, the midpoint of the lower edge of a box is selected as the contact point between the target and the ground, of which the pixel coordinates are used to calculate target depth. It is important to note that an accurate measurement of the height and the pitch angle for the camera are necessary to the P3D method. The overall algorithm is described in Algorithm 1.
@@
4:     if $u,v$ beyond the boundaries of $I$ then
@@
7:     $z\leftarrow f(K,H_{C},\alpha,\gamma)$ using Eq. 7
8:     $x\leftarrow(u-C_{u})\times z/f_{u}$ using Eq. 8
9:     $y\leftarrow(v-C_{v})\times z/f_{v}$ using Eq. 8
@@
As shown in Fig. 11(a), we depict the change in the $z$-component of the car’s coordinates against its longitudinal displacements for each experiment, with the score for the goodness of linear fit shown on the right side in (a), as an indicator of the ground flatness. Note that the detection range varies across the sampled cameras due to environmental occlusions and road corners, and the goodness of fit scores can be easily clustered into even, partially-even and uneven. Sample images of the experimental scenes can be found in Fig. 11(b) (g).

### v3

@@ 1 Introduction
Roadside perception (RP) technology is a key component in the intelligent transportation and autonomous driving domains, enabling features such as traffic flow management[45], decision support for autonomous vehicles[33], and campus surveillance and security[26], where the pursuit of reliable perception methods and high-quality datasets remain enduring topics.
@@
On a university campus, we have tested some open-sourced state-of-the-art algorithms in the roadside perception field such as BEVHeight[48], BEVDepth[20] and BEVFormer[22] with DAIR-V2X-I[53] and corresponding pre-trained models. When applied to campus roads characterized by complex living scenes and natural elements, these algorithms demonstrated unsatisfactory performance, failing to effectively distinguish between various objects and backgrounds, as illustrated in Fig.˜1. This phenomenon indicates that, despite their outstanding performance on urban traffic datasets like DAIR-V2X[53] and Rope3D[50], these algorithms still need improvement in terms of generalization capability and robustness when faced with more diverse environments.
@@
The highlighted observation points towards a research frontier in which residential areas akin to university campuses provide distinctive research vantages and real-world deployment scenarios, thereby broadening the scope of roadside perception technology development beyond the traditional confines of urban streets. Motivated by this, we have innovatively constructed a roadside perception dataset named CORP (Campus-Oriented Roadside Perception), which is the first large-scale, multi-modal roadside dataset specifically designed for campus and large public spaces. Within the CORP dataset, in response to the unconstrained nature of the campus environment, multiple cameras and lidars with diverse specifications have been employed and installed at varying heights and locations on roadside poles across the campus, capturing rich scene information from various perspectives. Table˜1 shows that according to the number of sensors and the complexity of scenes, we have partitioned the CORP dataset into three regions, namely Region A, Region B, and Region C, as shown in Fig.˜2.
@@ 2.1 Roadside Perception Datasets
Datasets exist for perception tasks in campus settings. The ShanghaiTech[25] dataset is designed for anomaly detection with images, which contains videos taken from multiple scenes in a university in Shanghai, and hundreds of footages are labeled with various types of abnormal events. The NWPU Campus[6] is built for the same task on a larger scale. The Usyd Campus[58] dataset provides a multi-modal collection over and around the University of Sydney campus for a continuous period of 1.5 years, focusing on vehicle-side perception. The Campus3D[19] dataset was generated through photogrammetric processing of UAV images taken over the National University of Singapore (NUS), which provides a dense point cloud of a campus tailored for point segmentation tasks. A comparison between CORP and the aforementioned datasets can be found in Table˜2.
@@ 3.1 Sensor Configuration
In the CORP dataset, the data frames were meticulously accumulated within a university campus, incorporating high-resolution imagery from 18 traffic surveillance cameras and spatial information captured by an ensemble of 9 automotive-grade LiDAR sensors. Each of these sensors was strategically installed atop intelligent poles at varying heights ranging from 2.5 to 6.5 meters above the ground, thereby ensuring comprehensive and diversified coverage for robust roadside perception tasks and presenting new challenges. An overview of the point cloud pattern and the position of cameras is shown in Fig.˜2 with a real-world picture of the pole. The range of the whole area extends about 700 meters in the west-east direction and has a north-south length of 350 meters. We divide the area into three regions, labeled A, B, and C, based on the detection range, and color-coded with blue, red, and black, respectively. We have configured different sensor layout schemes for different regions, and the specific number and distribution details of the sensors in the layout are shown in Table 1.
@@
We have constructed four types of coordinate systems for the CORP dataset, of which two are used for LiDAR sensors, one for cameras, and the last one for the global map, as illustrated in Fig.˜3. For LiDARs, the LiDAR-ego reference frame originates from the sensor itself with the $x$-axis pointing forward and $z$-axis upward. For the convenience of researchers and to safeguard data privacy, we have constructed an additional virtual LiDAR coordinate system, termed as LiDAR-base, transforms the origin of LiDAR-ego to the ground contact point of its host pole to comply its $x$-$o$-$y$ plane with the local ground surface. For cameras, the origin of the camera-ego coordinate system overlaps the optical center of the sensor with the $z$-axis pointing forward along the camera’s view and $x$-axis pointing leftward accordingly. The LiDAR and camera coordinate systems are employed to represent objects in the sensor’s field of view, while the map coordinates unify the objects into a global picture. The map coordinate system has an origin with known GPS, its $x$-axis coincides with the east and the $y$-axis with the north.
@@
Intrinsic parameters of a camera define the projection between the incoming rays and image pixels. For CORP, we use chessboards as explicit targets and employ MATLAB Camera Calibrator [11] following the method outlined by Zhang et al.[57] to determine the intrinsic calibration. This procedure yields a set of quantities including $f_{x},f_{y},c_{x},c_{y},k_{1},k_{2},p_{1},p_{2},k_{3}]$, where $f_{x}$ and $f_{y}$ denote the focal lengths, while $c_{x}$ and $c_{y}$ represent the principal points. Additionally, the parameters $k_{1}$, $k_{2}$, and $k_{3}$ pertain to radial distortion, and $p_{1}$ and $p_{2}$ are related to tangential distortion.
@@ 3.2 Data Acquisition and Annotation
Data Labeling and Visualization. Human annotators are employed for joint labeling of the 2D and 3D targets in the dataset using the format in DAIR-V2X, with additional information such as target ID and motion status (movingv.s. static) added. To generate pixel segmentation masks for the moving instances in mages, the foundation vision model SAM[16] is applied with the integration of human intervention and manual guidance. Visualization of data samples is shown in Fig.˜4. The images and point clouds can also be viewed with their annotations from SUSTechPoints[18] with our development toolkit.
@@ 3.4 Features and statistics
To get an overview of how targets are distributed in space, we aligned all identified 3D objects into their own LiDAR-base coordinate system, each of which has $x$-axis pointing eastward and $y$-axis towards the north, and visualized them collectively as shown in Fig.˜5(a), where warmer color indicates a higher concentration of targets. Also, it is not surprising that the spatial distribution of targets exhibits a discernible trend toward specific directions as also demonstrated in their orientation profile in Fig.˜5(c), since the mobility of targets is partially subjected to the road network in the environment. The allocation of targets indicates that our dataset encompasses plenty of road arrangements. Fig.˜5(b) showed the distribution profile for the distance between targets and their center LiDARs under BEV. The portion of targets decreases as the distance increases since point clouds are sparser there, but most of the targets appear relatively evenly within 100 meters away from the center.
@@
To glance at how targets distribute differently in CORP than in urban arterial streets, we analyzed the target densities of CORP, DAIR-V2X-i[53] and Rope3D[50]. As shown in Fig.˜6(a) and (c). CORP displays a sparsity of per-frame target numbers in both 2D and 3D cases, about half of the frames contain less than 15 targets while the other half has up to 50 targets per frame, which is not the case in its urban counterparts. However, the sparsity of targets does not imply simplicity of perception tasks, since accurately and efficiently localizing them from the background requires extra engineering for the model. In addition,
Although the volume of the object class is less than that of DAIR-V2X and A9, we found no significant imbalance in the number of targets, indicating a relatively minor disparity over categories, thus diminishing the risks of label inequity. Note that, Rope3D has no 2D bounding box annotations and therefore is excluded in Fig.˜6(c) and (d).
@@ 4.1 Baseline Methods
For LiDAR-based baseline, we trained PointPillars with a detection range customized to [-102.4, -102.4,-2.0, 102.4, 102.4, 4.0] and the size of the voxel grid as [0.16, 0.16] along $x$ and $y$ axes. All point clouds are transformed with annotations to the LiDAR-base coordinate system mentioned in Section˜3. For camera-based detection, we compared the performance of BEVHeight and ImVoxelNet. In addition, we trained BEVFusion following [24] but confined to perception sites with both camera and LiDAR installed. We adapt AB3D to the detection results as our tracking baseline.
@@ 4.2 Results
For 2D detection with CORP, we compared the performance of broadly acknowledged YOLO-V5 and its V8 counterparts. As shown in Table 4, the two methods demonstrate an obvious gap in overall precision, recall and AP scores. Despite the higher precision for YOLO-V5, V8 wins in terms of $AP_{50-95}$, implying a better generalization ability. However, challenges persist in accurately detecting distant targets, particularly when they are occluded by foreground elements or under suboptimal light conditions as shown in Fig.˜7(a).
@@
For the evaluation of P3D, we compared it with the traditional yet industrially-recognized IPM method[3] in terms of prediction error in the distance from the camera to the target. For a fair comparison, images from 18 cameras in CORP are examined within their corresponding region, categorized into A, B and C based on the detection range. An experiment vehicle equipped with RTK GPS is driven through the FOV of each camera slowly, then the distance from the vehicle to the camera is predicted by P3D and IPM independently. As shown in Fig.˜8, although for region A, the prediction error of P3D is larger in the middle but the overall error is less than 2 meters for that interval. Also we observed that P3D consistently outperformed IPM across all scenarios featuring diverse ground surfaces, camera height and pose angles, thereby establishing itself as a robust baseline for object-level distance estimation tasks. Again, here we focus on non-learning distance estimation methods.
@@ 5 Conclusion
In roadside scenarios, the camera sensor is usually positioned at a certain height $H_{c}$ above a local ground plane. We define the camera and the road coordinate systems as shown in Figure Fig.˜9. By employing P3D, the objective we aspire to achieve is to lift the pertinent 2D objects into real-world environment by the reconstruction of their 3D coordinates, in a non-learning fashion.
@@
To describe the ground plane in the camera’s frame, we use the transformation matrix $T_{road2cam}\in R^{4\times 4}$ to establish the relationship between the road and the camera coordinates. Any 3D point in the road system $x,y,z]_{road}^{T}$ can be transformed to the image plane to obtain their pixel coordinates $u,v]^{T}$ via the following two steps:
@@
where $Q_{cam}$ is the coordinates of a point in the camera system, $n=\begin{bmatrix}r_{21}&r_{22}&r_{23}\end{bmatrix}^{T}$ is the normal vector to the plane and $d$ is the intercept parameter, which is $H_{c}$ in the considered case. When it comes to a different plane, $d$, $\alpha$ and $\gamma$ should all be accounted for from that plane. By solving the above equations for $z$ in the case where $\gamma$ is close to 0, this depth $z$ can be approximated as:
@@
Once $z$ of a target in an image is obtained with its pixel coordinates $u,v]^{T}$, we can further get its Cartesian coordinates $x,y,z]_{cam}^{T}$ in the camera system by the following:
@@
Worth mentioning, the vanishing point of a camera, defined by the intersection of two mutually parallel lines in $o$-$x$-$y$-$z$ coordinate system, is considered to have infinite depth, i.e.the dominator in Eq.˜7 on the right side is zero, therefore we have:
@@
For the completeness of description, the overview of our Pseudo-3D (P3D) method for target distance estimation is shown in Fig.˜10. The processing pipeline of P3D is the following: (i) with a given image, perform line feature detection for parallel lines, which in most cases are lane markings, then find their intersections as the vanishing point. (ii) using the vanishing point to obtain the pitch angle of a camera based on Eq.˜9. (iii) with the given pixel coordinates $u,v]^{T}$ of a target, compute the estimated depth $z$ using Eq.˜7 and $x$, $y$ using Eq.˜8.
@@
Once the intrinsic parameters, pose angles and the height of camera sensors are measured, we can lift a 2D target in an image to a 3D point in the camera coordinate system by following the closed-form Eq.˜7 and Eq.˜8 with no computational cost, given an image-based 2D detector employed beforehand to produce the bounding boxes of interested targets. Subsequently, the midpoint of the lower edge of a box is selected as the contact point between the target and the ground, of which the pixel coordinates are used to calculate target depth. It is important to note that an accurate measurement of the height and the pitch angle for the camera are necessary to the P3D method. The overall algorithm is described in Algorithm˜1.
@@
4:  if $u,v$ beyond the boundaries of II then
@@
7:  $z\leftarrow f(K,H_{C},\alpha,\gamma)$ using Eq.˜7
8:  $x\leftarrow(u-C_{u})\times z/f_{u}$ using Eq.˜8
9:  $y\leftarrow(v-C_{v})\times z/f_{v}$ using Eq.˜8
@@
As shown in Fig.˜11(a), we depict the change in the $z$-component of the car’s coordinates against its longitudinal displacements for each experiment, with the score for the goodness of linear fit shown on the right side in (a), as an indicator of the ground flatness. Note that the detection range varies across the sampled cameras due to environmental occlusions and road corners, and the goodness of fit scores can be easily clustered into even, partially-even and uneven. Sample images of the experimental scenes can be found in Fig.˜11(b) (g).